Three consecutive terms of a geometric sequence are x - 3, 6 and x + 2.
Find the possible values of x.

Answer by Boreal(15235)

the ratios between adjacent products should be equal
6/(x-3)=(x+2)/6
cross multiply
x^2-x-6=36
x^2-x-42=0
(x-7)(x+6)=0
x=-6, 7 ANSWER
For -6, the series is -9,6,and -4, a ratio of -(2/3)
For 7, the series is 4,6,9 for a ratio of 2/3